OpenStudy (anonymous):

Please help, i give medals. solve for x: 3^x+4 = 6^2x-5.

OpenStudy (anonymous):

@HelpBlahBlahBlah @Collielver123

OpenStudy (helpblahblahblah):

i think x = 5.37374 not completly sure

OpenStudy (anonymous):

did you mean 3^(x+4)=6^(2x-5)

OpenStudy (anonymous):

yeah, that's what i meant

OpenStudy (helpblahblahblah):

lol ten the asnwer is x = 5.37374 (i looked it up :) )

OpenStudy (anonymous):

how did you do it?

OpenStudy (helpblahblahblah):

one sec

OpenStudy (helpblahblahblah):

http://answers.yahoo.com/question/index?qid=20110123150005AAWQMXs thats where i got the answer from

OpenStudy (anonymous):

thank you!

OpenStudy (helpblahblahblah):

anytime! lol